Find 4 consecutive odd integers where the product of the two smaller numbers is 64 less than the product of the two larger numbers.
Answer by stanbon(75887)

Find 4 consecutive odd integers where the product of the two smaller numbers is 64 less than the product of the two larger numbers.
----
1st: 2x-3
2nd: 2x-1
3rd: 2x+1
4th: 2x+3
------------------
Equation:
(2x-3)(2x-1) = (2x+1)(2x+3)-64
---
4x^2-8x+3 = 4x^2+8x+3 - 64
----
-16x = -64
x = 4
----
1st: 2x-3 = 5
2nd: 7
3rd: 9
4th: 11
--------------
Cheers,
Stan H.
============